OpenStudy (anonymous):

The volume of a cylinder is found using the formula , where r is the radius of the base and h is the height. The volume of a rectangular prism of the same width (2r) and height h, as the cylinder, is found by multiplying the area of the base by the height. What is the ratio of the volume of the cylinder to the volume of the prism?

OpenStudy (triciaal):

let Vc =volume of cylinder = pi*r^2*h let Vp = volume of prism = area of base * h area of base = l* w and w = 2r = l*2r*h ratio of Vc to Vp = (pi*r^2*h)/(l*2r*h) pi*r/2l hope this helps